OpenStudy (anonymous):

Solve for x: (2x - 4)^2 = 1 Can someone please just refresh my memory on how to solve something like this, completely forgot how you would.

OpenStudy (anonymous):

something, when squared, is equal to 1 (-1)^2 = 1 and (1)^2 = 1 therefore, the innards must equal -1 or 1 2x - 4 = 1 or 2x - 4 = -1

OpenStudy (anonymous):

Thank You!(:

OpenStudy (anonymous):

youre welcome